Your annual salary at the beginning of the year was $53,000. During the year you received raises totaling $1,500. During the same year, the rate of inflation was 2.9%. Taking into account your raise and the rate of inflation, how much did your purchasing power change during the year?

Enter your answer as a percentage rounded to two decimal places (e.g., enter -1.34%, NOT -0.0134).
Answer #1

Answer:

Initial Salary S=$53000

Salary hike H=$1500

So Rise in Salary in % R=1500/53000=2.83%

Inflation rate I=2.9%

So change in purchasing power=R-I=2.83%-2.9%=-0.07%